# Graph Neural Network Enhanced Retrieval for Question Answering of LLMs

### v1

@@ 1 Introduction
Large language models (LLMs) continue to struggle with factual errors when encountering knowledge intensive questions [7, 19, 9]. Although retrieval-augmented LLMs [16, 15] have improved factuality and precision of question answering by including relevant passages, there remains a persistent challenge in accurately capturing all the relevant supporting passages when encountering complex knowledge-intensive questions. This limitation can be attributed to the inherent information asymmetry in complex questions. In particular, the questions tend to consist of elaborate background details, leaving only a small portion dedicated to specific inquiries. As an example, in question: ‘Why did crime rise on Mars after the Mafia’s arrival?’, the majority part delves into the background (i.e., ‘crime rise …… Mafia’s arrival’) with only a few words requesting the reason (i.e., ‘Why’), which consequently retrieves passages that focus on the details of crime rising instead of the reason for it.
This phenomenon also frequently arises in multi-hop reasoning questions. Considering the sample question ‘Where was the performer of song Left & Right (Song) born?’ in Fig. 1, it becomes apparent that while we can retrieve the knowledge that the performer of song Left & Right is D’Angelo, his birthplace remains absent. Previous works [30, 23, 35] have attempted to address this issue by incorporating multi-hop reasoning or question rewriting into retrieval processes, which enables them to retrieve information based on prior reasoning outcomes. LLMs, however, often generate plausible reasons and incorrect rewriting questions without accurate prior domain knowledge of the given question [7, 36], thus affecting the subsequent retrieval process.
@@
One reason why existing methods struggle to handle the information asymmetry is their tendency to consider passages in isolation [14] and retrieve the supporting passages based on semantic distances, making it difficult to retrieve all the supporting passages, especially those containing only few words for inquiry. However, the supporting passages of background and inquiry information are usually correlated. They can be structure-related when located in the same section or document, e.g., the happened event (i.e., crime rises on Mars) and its corresponding reason are located in the same section but different passages. Also, they can be keyword-related by sharing the same keywords or entities, e.g., the passages of background information (performer of song Left &\&& Right) and inquiry information (birthplace of D’Angelo) in Fig. 1 share the same keyword: D’Angelo. By considering the relatedness between passages, it is possible to retrieve all the supporting passages even when they have significant semantic differences from the query. In this work, we aim to enhance retrieval by taking the relatedness between passages into account.
@@
Contributions. To establish the relatedness between passages for retrieval purposes, this study initially constructs a graph of passages by connecting individual passages based on both structural information and shared keywords, with each passage as a node in this graph. The key challenge lies in how to effectively leverage the passage of graphs to enhance retrieval coverage. Graph neural networks (GNNs) are neural networks tailored to analyze graph data and adeptly grasp the relationships between nodes and edges [26]. Thus, we propose to leverage a GNN to enhance the retrieval process by effectively capturing the relatedness between passages and name this method GNN-Ret. The GNN facilitates the integration of semantic distances between related passages, thus enabling the semantic distances of passages containing background information to impact the retrieval of passages relevant to the inquiry.
@@
Figure 2: Accuracy and average query times of LLMs for our proposed methods and baselines on 2WikiMQA dataset.
To address multi-hop reasoning questions, we further propose a retrieval method, named RGNN-Ret, which leverages a recurrent graph neural network (RGNN) to enhance the retrieval process at each step by integrating the retrievals from previous steps through the interconnected graphs of passages. This boosts the retrieval coverage for supporting passages over steps even when LLMs generate incorrect reasons or subquestions for retrieval.
@@
Through the experiments on four benchmark datasets, we demonstrate the effectiveness of GNN in enhancing retrieval coverage for supporting facts and thus improving accuracy for QA tasks. For example, as shown in Fig. 2, our proposed GNN-Ret with a single query of LLMs significantly outperforms baselines in accuracy on the 2WikiMQA dataset [6], including those methods that require multiple queries of LLMs, which demonstrates the effectiveness of GNNs in improving retrieval coverage of supporting passages. Moreover, by extending GNN to multi-hop reasoning questions, our proposed RGNN-Ret achieves state-of-the-art accuracy.
@@ 2 Related Works
Retrieval-augmented LLM. QA [31] is a task that often requires external and up-to-date knowledge sources to answer factoid-based questions. Retrieve-and-read is the basic framework for these questions [5, 37]. For retrieval, sparse retrievers, e.g., TF-IDF and BM25 [24], or dense retrievers, e.g., DPR [13] and Contriever [8], are used to compute lexical distances or semantic distances between passages and the question. The passages with smaller distances are retrieved, which are then prefixed with the question for factual answering [16]. However, this framework treats passages as isolated units, making it difficult to retrieve all the supporting passages for the question in a single step. To address this limitation, many approaches have been proposed to integrate the retrieval and reasoning processes to improve the retrieval of supporting passages [5, 27, 3, 30, 11, 35, 23]. They first prompt LLMs to generate the next-step reason or subquestion and then use it to guide retrieval or knowledge search for QA. However, a challenge in these methodologies pertains to the potential generation of hallucinated reasons or erroneous subquestions by LLMs without accurate domain knowledge of this question, leading to degraded subsequent retrievals [7].
Graph-enhanced LLM. As structured, explicit, and editable representations of knowledge, knowledge graphs (KGs) have been effectively used for boosting retrieval coverage [20] and enhancing reasoning capabilities of LLMs [17, 34, 1, 32, 22, 2]. Instead of naively combining KGs with LLMs, many works directly prompt LLMs to perform reasoning on KGs [28, 10, 18]. They first identified the entities on the knowledge graph and then search reasoning paths accordingly [28, 18, 10]. However, KG-based methods encounter a limitation that they can only handle questions that are effectively represented as KGs. They may struggle with questions that necessitate a comprehensive understanding of lengthy contextual information, such as the mentioned example: the cause of crime increasing. Rather than KGs, a recent study [12] attempts to construct a graph of passages and perform reasoning on this graph. It connects documents according to some specific words and performs reasoning on this graph for QA. In contrast, we utilize the graph of passages to enhance the retrieval process. Another related work, KGP [33], performs retrieval on the graph and search related passages. However, this approach does not consider the integration between related passages and may not effectively enhance retrieval coverage for the supporting passages with poor semantic similarity.
@@ 3 Graph Neural Network Enhanced Retrieval for QA
In this section, we aim to enhance the retrieval coverage of supporting passages by leveraging the relatedness between passages for QA tasks. Specifically, we construct a graph of passages by connecting passages that are structural-related and keyword-related, where each node represents a passage (Section 3.2). To exploit the relatedness between passages to enhance retrieval, we utilize a GNN to enable integration between related passages (Section 3.3). For multi-hop questions, we further leverage an RGNN to establish relationship between graphs over steps to enhance retrieval for the multi-hop answering process (Section 3.4).
3.1 Problem Description
The QA system typically adapts a retrieve-and-read routine [16]: the retriever computes semantic distances $D_{q}$ between passages and question $q$ in the embedding space, finds out the relevant passages $P_{q}$ from the whole corpus $P$, and then the reader (LLM) generates the answer $a$ based on them. This process can be formulated as follows:
$D_{q}=\{\phi(q,p)|p\in P\},\quad P_{q}=\text{Retrieve}(q,D_{q}),\quad a=\text{Read}(q,P_{q}),$ (1)
where $\phi(q,p)$ measures the semantic distance between question $q$ and passage $p$. A smaller value of $\phi(q,p)$ indicates higher similarity between $q$ and $p$. Considering the information asymmetry of complex questions, directly using the initial semantic distances of passages $D_{p}$ for retrieval is difficult to retrieve all the supporting passages. Therefore, we propose to process $D_{q}$ by taking the relatedness between passages into account and thus enhance the retrieval for QA.
3.2 Graph of Passages
@@
Establish relationships between related passages. Existing retrieval methods perform retrieval merely relying on semantic distances between passages and the question [5, 37], which may fail to retrieve some supporting passages with large semantic distances to the question due to the information asymmetry. However, these supporting passages are usually related to those with small semantic distances to the question. They can be structural-related, e.g., the happened event (‘crime rising’) and its corresponding reasons are located in the same section. As shown in Fig. 1, they can also be keyword-related, e.g., the supporting passages for background information (‘performer of song Left &\&& Right’) and inquiry information (‘the birthplace of D’Angelo’) share the same keyword (‘D’Angelo’). Therefore, we establish relationship between passages using structural information and shared keywords to construct the graph of passages. Specifically, when chunking the documents, we record the order of passages and connect passages that are physically next to each other in the documents. In addition, we extract keywords from passages by prompting LLMs and connect passages containing the same keywords [20, 33]. As such, the related passages are connected, having the potential to facilitate retrieval.
3.3 GNN Enhanced Retrieval on Graph of Passages
@@
After establishing connections between related passages, we obtain a graph of passages, with each passage as a node. When conducting the retrieval process, we first compute the semantic distances between passages and the question $q$ [5, 37]. The semantic distances of related passages are connected according to the graph of passages. To take relatedness between passages into account, we utilize a GNN to process semantic distances based on the graph of passages and obtain the integrated semantic distances, which benefits the retrieval of supporting passages.
@@
Graph neural network. A graph is defined as an ordered pair $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{G}$ is the set of nodes $v_{i}$ and $\mathcal{E}$ is the set of edges $e_{ij}$ connecting two nodes $v_{i}$ and $v_{j}$. We define $h_{i}^{l}$ as the hidden state (integrated semantic distance) of the $i-$th node for $i\in\{1,\cdots,|\mathcal{V}|\}$ at layer $l$. We input semantic distances of passages into GNN as $h_{i}^{0}$. The set of nodes $\mathcal{N}_{v_{i}}$ stands for the neighbors of node $v_{i}$. For an $L-$layer GNN [26], the hidden state of each node $v_{i}\in\mathcal{V}$ is updated through iterative interactions with neighbors. Specifically, at layer $l$, the received messages $m_{i}^{l}$ of node $v_{i}$ are calculated using the hidden states of its neighbors. Then the received messages $m_{i}^{l}$ and the previous-layer hidden state $h_{i}^{l-1}$ are utilized to compute the hidden state $h_{i}^{l}$ at layer $l$, which are formulated as follows:
$m_{i}^{l}=f_{m}^{l}(h_{j}^{l-1}|\forall j\in\{1,\cdots,|\mathcal{N}_{v_{i}}|\}),\quad h_{i}^{l}=f_{h}^{l}(m_{i}^{l},h_{i}^{l-1}),$ (2)
where functions $f_{m}^{l}(\cdot)$ and $f_{h}^{l}$ aggregate neighbors’ information and update the hidden state, respectively. By processing hidden states for all the passages layer by layer, we obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}\in\mathcal{V}$. We elaborately introduce the GNN for retrieval as below.
@@
Minimum semantic distance as message. Since each passage maintains and shares many keywords with other passages, each passage has a large number of neighbors. Some of them are not relevant to the question, contributing to distracting or misleading information. Therefore, each node $v_{i}$ only receives the message from the neighbor that has the minimum (integrated) semantic distance to the question at layer $l$. The received message of node $v_{i}$ is thus formulated as $m_{i}^{l}=\min_{j\in\mathcal{N}_{v_{i}}}h_{j}^{l-1}$.
@@
Relevant nodes sampling. Given the large scale of the graph constructed by plenty of passages, it is computationally inefficient to propagate information across all nodes. Besides, allowing irrelevant nodes to pass messages to their neighbors may affect the retrieval process. Therefore, we only sample the relevant node set $\mathcal{T}_{K}^{l}$ with the top $K$ smallest (integrated) semantic distances at layer $l$. Only neighbors of relevant nodes receive messages from them and update hidden states. We define $\mathcal{S}_{K}^{l}$ as the neighbor set of the relevant node set $\mathcal{T}_{K}^{l}$. For each node $v_{i}\in\mathcal{S}_{K}^{l}$, the message passed from neighbors that are relevant nodes is reformulated as follows:
@@
The nodes without received message maintain their hidden states of the previous layer. For the GNN with $L$ layers, we compute the hidden state layer by layer and then obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}$. With GNN, the supporting passages for inquiry information integrate with the small semantic distance from those for background information and thereby obtain smaller integrated semantic distances, promoting them to be retrieved.
Figure 3: Illustration of GNN-Ret and RGNN-Ret.

### v2

# Graph Neural Network Enhanced Retrieval for Question Answering of Large Language Models

@@ Abstract
Graph Neural Network Enhanced Retrieval for Question Answering of Large Language Models
Zijian Li12thanks: Work done during an internship at Microsoft Research Asia.   , Qingyan Guo23, Jiawei Shao1, Lei Song2, Jiang Bian2, Jun Zhang1thanks: Corresponding Author.  , Rui Wang2† 1Hong Kong University of Science and Technology 2Microsoft Research 3Tsinghua University {zijian.li,jiawei.shao}@connect.ust.hk, gqy22@mails.tsinghua.edu.cn, {lesong,jiabia,ruiwa}@microsoft.com, eejzhang@ust.hk
@@ 1 Introduction
Large language models (LLMs) continue to struggle with factual errors when encountering knowledge intensive questions (Huang et al., 2023; Mallen et al., 2022; Ji et al., 2023). Although retrieval-augmented LLMs (Lewis et al., 2020) have improved factuality and precision of question answering by including relevant passages, there remains a persistent challenge in accurately capturing all the supporting passages when encountering complex knowledge-intensive questions. This limitation can be attributed to the inherent information asymmetry in complex questions. In particular, the questions tend to consist of elaborated background details, leaving only a small portion dedicated to specific inquiries. As an example, in question: ‘Why did crime rise on Mars after the Mafia’s arrival?’, the majority part delves into the background (i.e., ‘crime rise …… Mafia’s arrival’) with only a few words requesting the reason (i.e., ‘Why’), which consequently retrieves passages on details of crime rising instead of the reason for it.
This phenomenon also frequently arises in multi-hop reasoning questions. Considering the sample question ‘Where was the performer of song Left & Right (Song) born?’ in Fig. 1, it becomes apparent that while we can retrieve the knowledge that the performer of song Left & Right is D’Angelo, his birthplace remains absent. Previous works Trivedi et al. (2023); Press et al. (2022); Yao et al. (2022) have attempted to address this issue by incorporating multi-hop reasoning or question rewriting into retrieval processes, which enables them to retrieve information based on prior reasoning outcomes. LLMs, however, often generate plausible reasons and incorrect rewriting questions without accurate prior domain knowledge of the given question Huang et al. (2023); Zhang et al. (2023), thus affecting the subsequent retrieval process.
@@
One reason why existing methods struggle to handle the information asymmetry is their tendency to consider passages in isolation Karpukhin et al. (2020a) and retrieve the supporting passages based mainly on semantic distances, making it difficult to retrieve all the supporting passages, especially those containing only few words for inquiry. However, the supporting passages of background and inquiry information are usually correlated. They can be structure-related when located in the same section or document, e.g., the happened event (i.e., crime rises on Mars) and its corresponding reason are located in the same section but different passages. Also, they can be keyword-related by sharing the same keywords or entities, e.g., the passages of background information (performer of song Left &\&& Right) and inquiry information (birthplace of D’Angelo) in Fig. 1 share the same keyword: D’Angelo. By considering the relatedness between passages, it is possible to retrieve all the supporting passages even when they have significant semantic differences from the query. In this work, we aim to enhance retrieval by taking the relatedness between passages into account.
@@
Contributions. To establish the relatedness between passages for retrieval purposes, this study initially constructs a graph of passages by connecting individual passages based on both structural information and shared keywords, with each passage as a node in this graph. The key challenge lies in how to effectively leverage the passage of graphs to enhance retrieval coverage. Graph neural networks (GNNs) are neural networks tailored to analyze graph data and adeptly grasp the relationships between nodes and edges Scarselli et al. (2008). Thus, we propose to leverage a GNN to enhance the retrieval process by effectively capturing the relatedness between passages and name this method GNN-Ret. The GNN facilitates the integration of semantic distances between related passages, thus enabling the semantic distances of passages containing background information to impact the retrieval of passages relevant to the inquiry. To address multi-hop reasoning questions, we further propose a retrieval method, named RGNN-Ret, which leverages a recurrent graph neural network (RGNN) to enhance the retrieval process at each step by integrating the retrievals from previous steps through the interconnected graphs of passages. This boosts the retrieval coverage for supporting passages over steps even when LLMs generate incorrect reasons or subquestions for retrieval.
@@
Figure 2: Accuracy and average number of LLM queries for our proposed methods and baselines on 2WikiMQA.
@@
Through the experiments on four benchmark datasets, we demonstrate the effectiveness of GNNs in enhancing retrieval of supporting passages and thus improving accuracy for QA. For example, as shown in Fig. 2, our proposed GNN-Ret with a single query of LLMs significantly outperforms baselines in accuracy on 2WikiMQA Ho et al. (2020), including those methods that require multiple queries of LLMs. Moreover, by extending GNN to multi-hop reasoning questions, our proposed RGNN-Ret achieves state-of-the-art accuracy.
@@ 2 Related Works
Retrieval-augmented LLM. QA Voorhees et al. (1999) is a task that often requires external and up-to-date knowledge sources to answer factoid-based questions. Retrieve-and-read is the basic framework for these questions Gao et al. (2023); Zhu et al. (2023). For retrieval, sparse retrievers, e.g., TF-IDF and BM25 Robertson et al. (2009), or dense retrievers, e.g., DPR Karpukhin et al. (2020b) and Contriever Izacard et al. (2021), are applied to compute lexical distances or semantic distances between passages and the question. The passages with smaller distances are retrieved, which are then prefixed with the question for factual answering Lewis et al. (2020). However, this framework treats passages as isolated units, making it difficult to retrieve all the supporting passages for the question in a single step. To address this limitation, many approaches have been proposed to integrate the retrieval and reasoning processes to improve the retrieval of supporting passages Gao et al. (2023); Shao et al. (2023); Feng et al. (2024); Trivedi et al. (2023); Jiang et al. (2023b); Yao et al. (2022); Press et al. (2022). They first prompt LLMs to generate the next-step reason or subquestion and then use it to guide retrieval for QA. However, a challenge in these methodologies pertains to the potential generation of hallucinated reasons or erroneous subquestions by LLMs without accurate domain knowledge of this question, leading to degraded subsequent retrievals Huang et al. (2023).
Graph-enhanced LLM. As structured, explicit, and editable representations of knowledge, knowledge graphs (KGs) have been effectively used for boosting retrieval coverage Min et al. (2019) and enhancing reasoning capabilities of LLMs Li et al. (2023a); Xie et al. (2022); Baek et al. (2023); Wang et al. (2023); Park et al. (2023); Ding et al. (2024); Jin et al. (2023); Li et al. (2023b); Xiong et al. (2024); He et al. (2024). Instead of naively combining KGs with LLMs, many works directly prompt LLMs to perform reasoning on KGs Sun et al. (2024); Jiang et al. (2023a); LUO et al. (2024). They first identified the entities on the knowledge graph and then search reasoning paths accordingly Sun et al. (2024); LUO et al. (2024); Jiang et al. (2023a). However, KG-based methods encounter a limitation that they can only handle questions that are effectively represented as KGs. They may struggle with questions that necessitate a comprehensive understanding of lengthy contextual information, such as the mentioned example: the cause of crime increasing. Rather than KGs, a recent study Jin et al. (2024) attempts to construct a graph of passages and perform reasoning on this graph. It connects documents according to some specific words and performs reasoning on this graph for QA. In contrast, we utilize the graph of passages to enhance the retrieval process. Another related work, KGP Wang et al. (2024), performs retrieval on the graph and search related passages. However, this approach does not consider the integration between related passages and may not effectively enhance retrieval for the supporting passages with poor semantic similarity.
@@ 3 Graph Neural Network Enhanced Retrieval for QA
The QA system typically adopts a retrieve-and-read routine Lewis et al. (2020): the retriever computes semantic distances between passages and the question in the embedding space, finds out relevant passages from the whole corpus, and then the reader (LLM) generates the answer based on them. Considering the information asymmetry of complex questions, directly using the initial semantic distances of passages for retrieval is difficult to retrieve all the supporting passages. Therefore, we propose to process by taking the relatedness between passages into account. Specifically, we construct a graph of passages (GoPs) by connecting related passages (Section 3.1). To exploit the relatedness between passages to enhance retrieval, we utilize a GNN to enable integration between related passages (Section 3.2). For multi-hop questions, we further leverage an RGNN to establish relationship between graphs over steps to enhance retrieval over answering processes (Section 3.3).
3.1 Graph of Passages (GoPs)
@@
Establish relationships between related passages. Existing retrieval methods rely on semantic distances between passages and the question Gao et al. (2023); Zhu et al. (2023), which may fail to retrieve some supporting passages with large semantic distances, but related to those with small semantic distances to the question, due to the information asymmetry. These passages can be structural-related, e.g., the happened event (‘crime rising’) and its corresponding reasons are located in the same section. As shown in Fig. 1, they can also be keyword-related, e.g., the supporting passages for background information (‘performer of song Left &\&& Right’) and inquiry information (‘the birthplace of D’Angelo’) share the same keyword (‘D’Angelo’). Therefore, we propose to establish relationship between passages using structural information and shared keywords to construct the graph of passages. Specifically, when chunking the documents, we record the order of passages and connect passages that are physically next to each other in documents. In addition, we extract keywords from passages by prompting LLMs and connect passages containing the same keywords Min et al. (2019); Wang et al. (2024). As such, the related passages are connected, having the potential to facilitate retrieval.
3.2 GNN Enhanced Retrieval on GoPs
@@
After establishing connections between related passages, we obtain a GoPs, with each passage as a node. When conducting the retrieval process, we first compute the semantic distances between passages and the question $q$ Gao et al. (2023); Zhu et al. (2023). The semantic distances of related passages are connected according to the GoPs. To take relatedness between passages into account, we utilize a GNN to process semantic distances based on GoPs and obtain the integrated semantic distances, which benefits the retrieval of supporting passages.
@@
Graph neural network. A graph is defined as an ordered pair $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ is the set of nodes $v_{i}$ and $\mathcal{E}$ is the set of edges $e_{ij}$ connecting two nodes $v_{i}$ and $v_{j}$. We define $h_{i}^{l}$ as the hidden state (integrated semantic distance) of the $i-$th node for $i\in\{1,\cdots,|\mathcal{V}|\}$ at layer $l$. We input semantic distances of passages into GNN as $h_{i}^{0}$. The set of nodes $\mathcal{N}_{v_{i}}$ stands for the neighbors of node $v_{i}$. For an $L-$layer GNN Scarselli et al. (2008), the hidden state of each node $v_{i}\in\mathcal{V}$ is updated through iterative interactions with neighbors. Specifically, at layer $l$, the received messages $m_{i}^{l}$ of node $v_{i}$ are calculated using the hidden states of its neighbors. Then the received messages $m_{i}^{l}$ and the previous-layer hidden state $h_{i}^{l-1}$ are utilized to compute the hidden state $h_{i}^{l}$ at layer $l$. By processing hidden states for all the passages layer by layer, we obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}\in\mathcal{V}$. We elaborately introduce GNNs for retrieval as below.
@@
Minimum semantic distance as message. Since each passage maintains and shares many keywords with other passages, each passage has a large number of neighbors. Some of them are not relevant to the question, contributing to large semantic distances and misleading information. Therefore, each node $v_{i}$ only receives the message from the neighbor that has the minimum (integrated) semantic distance to the question at layer $l$. The received message of node $v_{i}$ is thus formulated as $m_{i}^{l}=\min_{j\in\mathcal{N}_{v_{i}}}h_{j}^{l-1}$.
Figure 3: Illustration of GNN-Ret and RGNN-Ret.
@@
Relevant nodes sampling. Given the large scale of GoPs with plenty of nodes, it is computationally inefficient to propagate information across all of them. Besides, allowing irrelevant nodes to pass messages to their neighbors may affect the retrieval process. Therefore, we only sample the relevant node set $\mathcal{T}_{K}^{l}$ with the top $K$ smallest (integrated) semantic distances at layer $l$. Only neighbors of relevant nodes receive messages from them and update hidden states. We define $\mathcal{S}_{K}^{l}$ as the neighbor set of the relevant node set $\mathcal{T}_{K}^{l}$. For each node $v_{i}\in\mathcal{S}_{K}^{l}$, the message passed from neighbors that are relevant nodes is reformulated as follows:
@@
The nodes without received messages maintain their hidden states of the previous layer. For the GNN with $L$ layers, we compute the hidden state layer by layer and then obtain the integrated semantic distance $h_{i}^{L}$ for each node $v_{i}$. With GNN, the supporting passages for inquiry information integrate with the small semantic distance from those for background information and thereby obtain smaller integrated semantic distances, promoting them to be retrieved.